OpenStudy (anonymous):

You have relatives living in the United Kingdom and in France. Suppose that you have purchased a prepaid phone card with a value of $75. Calls to the United Kingdom cost 23¢ per minute, while calls to France cost 21¢ per minute.

OpenStudy (anonymous):

(b) Graph the inequality, and discuss a possible solution in the context of the real-life situation

OpenStudy (anonymous):

I know the inequality is 0.21y+0.23x<=75... but how do I graph this? @ranga

OpenStudy (ranga):

Replace the inequality with equality and graph the boundary line. Then identify on which side of the boundary line the solution set lies and shade that region. And since this inequality involves a less than or EQUAL sign, the graph has to be a solid or bold or continuous line. If the inequality were a less than or greater than without the equal sign, then the boundary line should be a dotted line.

OpenStudy (ranga):

Boundary line: 0.21y + 0.23x = 75 0.21y = -0.23x + 75 y = -0.23/0.21*x + 75/0.21 y = -1.095x + 357.143

OpenStudy (anonymous):

What about the real life situation?

OpenStudy (ranga):

For real life situation you can say if you spend a lot of minutes for calls to the UK that would limit the number of minutes available for calls to France and vice-versa. So you can set a limit, say, a maximum of 100 minutes for calls to France and then using the equation find the maximum minutes that you can call UK.